Session: Feb 2000 (session 1)
Wavelength: 18 cm
To give feedback on this experiment, click
Medicina (C. Bortolotti at Fri Feb 18 19:15:59 MET 2000) Success
No known problems
Effelsberg (A. Kraus at Mon Feb 21 08:16:41 MET 2000) Minor failures
Due to snowfall the observation could not start before UT0955.
Torun (K.Borkowski at Mon Feb 21 09:26:08 MET 2000) Success
No known problems.
Noto (C.Trigilio at Mon Feb 21 12:10:20 MET 2000) Minor failures
Two scan lost from 5:30 to 5:43 on 3c286 for cable wrap problems.
Onsala (A. Polatidis at Mon Feb 21 12:49:41 MET 2000) Success
Westerbork (Tony Foley at Mon Feb 21 15:58:53 MET 2000) Success
No known problems. Full tied array used.
Jodrell (Alastair Gunn at Wed Feb 23 14:37:03 MET 2000) Minor failures
Some data missing because PI did not allocated sufficient
telescope slew times between target and calibrator. No
other known problems.
Calibration (Alastair Gunn at Wed Apr 19 13:03:03 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/ek010b.antab for details
Calibration (Alastair Gunn at Wed Apr 19 13:03:29 METDST 2000) Success
Calibration data available from vlbeer: see file feb00/eh007b.antab for details